# SafeMate: A Model Context Protocol-Based Multimodal Agent for Emergency Preparedness

### v1

@@ Abstract
Despite the abundance of public safety documents and emergency protocols, most individuals remain ill-equipped to interpret and act on such information during crises. Traditional emergency decision support systems (EDSS) are designed for professionals and rely heavily on static documents like PDFs or SOPs, which are difficult for non-experts to navigate under stress. This gap between institutional knowledge and public accessibility poses a critical barrier to effective emergency preparedness and response.
We introduce SafeMate, a retrieval-augmented AI assistant that delivers accurate, context-aware guidance to general users in both preparedness and active emergency scenarios. Built on the Model Context Protocol (MCP), SafeMate dynamically routes user queries to tools for document retrieval, checklist generation, and structured summarization. It uses FAISS with cosine similarity to identify relevant content from trusted sources such as CDC, FEMA, and SOHA.
@@
**footnotetext: These authors contributed equally to this work
@@ 1 Introduction
To address these limitations, we introduce SafeMate, a retrieval-augmented, multimodal AI assistant designed to enhance both emergency preparedness and response. Unlike conventional EDSS or static checklists, SafeMate delivers adaptive, verified, and structured outputs using advanced reasoning and knowledge retrieval methods.
SafeMate is built on a modular agent architecture enabled by the Model Context Protocol (MCP)(Anthropic, 2024). MCP provides a framework for orchestrating multiple external tools and services—such as document retrievers, geospatial databases, public preparedness knowledge bases, and search APIs, under a shared semantic interface. This architecture allows the agent to dynamically route queries to the appropriate tool, retrieve domain-specific information, and generate a unified, context-sensitive response. Unlike monolithic pipelines, this modularity supports rapid integration of new knowledge sources and tools from diverse domains, ranging from legal preparedness guidelines to localized infrastructure plans.
@@ 2.4 Cross-Domain Applications of Large Language Models
Across these domains, one clear insight emerges: LLM-based systems achieve greater reliability and impact when coupled with structured reasoning, external verification, and modular extensibility—core tenets embodied by SafeMate’s design.
@@ 3.1 System Architecture Overview
As illustrated in Figure 1, the SafeMate system is organized into a modular architecture centered around an LLM and orchestrated via the Model Context Protocol (MCP). The MCP Client layer, implemented using LangChain and LangGraph, manages interaction between the language model and external tools. It invokes FAISS-based retrieval over an embedded knowledge base for static information and routes dynamic tool calls—such as YouTube search, Google Maps rendering, or weather APIs—through LangChain MCP adapters. These tools are accessed via the MCP Server infrastructure, such as Smithery. A FastAPI backend exposes the LLM agent to users and facilitates requests via both web and geospatial interfaces. The response pipeline culminates in the LLM composing a multimodal, context-aware output, grounded in retrieved documents and enriched with tool-based augmentations.
@@ 3.3 Low-Hallucination Language Model: o3-mini
In emergency response applications, the accuracy of information is paramount. To minimize the risk of generating incorrect or misleading content, we selected OpenAI’s o3-mini language model for SafeMate. According to evaluations using Vectara’s Hughes Hallucination Evaluation Model, o3-mini exhibited a hallucination rate of 0.8%, outperforming other models such as 4o (1.4%) and DeepSeek r1 (14.3%)
@@ 3.6 Hierarchical Retrieval with RAPTOR
To effectively manage and retrieve information from extensive and complex emergency documents, SafeMate integrates the RAPTOR framework. RAPTOR constructs a hierarchical tree structure through recursive embedding, clustering, and summarization of text chunks, enabling multi-level abstraction and efficient information retrieval.
@@ 3.6.2 Dimensionality Reduction and Clustering
$p_{ij}=\exp\left(-\frac{\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}}{\sigma_{i}}\right)$
@@
$q_{ij}=\left(1+a\|\mathbf{y}_{i}-\mathbf{y}_{j}\|^{2b}\right)^{-1}$
@@
$\mathcal{L}_{\text{UMAP}}=\sum_{i\neq j}\left(p_{ij}\log\frac{p_{ij}}{q_{ij}}+(1-p_{ij})\log\frac{1-p_{ij}}{1-q_{ij}}\right)$
@@
$p(\mathbf{x}_{i})=\sum_{k=1}^{K}\pi_{k}\,\mathcal{N}(\mathbf{x}_{i}\mid\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k})$
@@
$\text{BIC}=\ln(N)\cdot k-2\cdot\ln(\hat{L})$
@@ 4.1 Evaluation and Verification
4.1.1 Evaluation Setup
Figure 3 illustrates our evaluation framework, where GPT serves as an automated evaluator to assess the quality of SafeMate’s generated responses. Given an emergency context that includes structured meta-information (e.g., disaster type, time, location) and supporting documents, a generator module synthesizes a question and a corresponding answer. SafeMate then generates its own response to the same question. Both responses are passed to an LLM-based evaluator (e.g., GPT-4), which scores the outputs across five dimensions: correctness, groundedness, completeness, relevance, and fluency. This evaluation loop enables scalable, consistent, and multi-criteria assessment of SafeMate’s response quality, particularly in settings where human evaluation is costly or infeasible.
@@ 4.2 Analysis
Groundedness and Hallucination Resistance.
@@ 5 Limitations and Future Work
SafeMate has primarily been tested using U.S.-centric emergency guidelines (FEMA, CDC). This restricts its applicability to other regions unless localized knowledge bases are incorporated. In future work, we aim to support multilingual KBs and culturally adaptive responses. Although retrieval accuracy and grounding have been evaluated, large-scale user testing in real disaster drills or field conditions remains future work. We plan to conduct scenario-based evaluations in collaboration with municipal emergency offices to quantify impact on decision-making speed and public trust.
@@ 6 Ethical Considerations
First, while SafeMate incorporates hallucination mitigation strategies—including the use of the o3-mini model and double-checking with retrieved evidence—no generative model is entirely free from factual errors. Misinterpretations of ambiguous queries, reliance on outdated documents, or edge-case generation failures could lead to inappropriate recommendations. To minimize potential harm, all outputs are explicitly labeled as informational and not legally or medically prescriptive.
@@ 7 Conclusion
We presented SafeMate, a retrieval-augmented, tool-integrated AI assistant designed to make emergency preparedness and response information more accessible, reliable, and actionable for the general public. Addressing the limitations of traditional emergency systems—namely their reliance on static documents, professional knowledge, and poor real-time usability—SafeMate integrates modular components including the Model Context Protocol (MCP), FAISS-based similarity retrieval, RAPTOR-style hierarchical summarization, and low-hallucination LLMs.

### v2

# SafeMate: A Modular RAG-Based Agent for Context-Aware Emergency Guidance

@@ Abstract
Despite the abundance of public safety documents and emergency protocols, many individuals remain inadequately prepared to effectively interpret and act upon such information during crises. Traditional emergency decision support systems (EDSS) are designed for professionals and rely heavily on static documents like PDFs or SOPs, which are difficult for non-experts to navigate under stress. This gap between institutional knowledge and public accessibility poses a critical barrier to effective emergency preparedness and response. To address this challenge, we introduce SafeMate, a retrieval-augmented AI assistant that delivers accurate, context-aware guidance to general users in both preparedness and active emergency scenarios. Built on the Model Context Protocol (MCP), SafeMate dynamically routes user queries to tools for document retrieval, checklist generation, and structured summarization. It uses FAISS with cosine similarity to identify relevant content from trusted sources. Experiments on emergency preparedness queries show that the proposed SafeMate outperforms GPT-4o and GPT-3.5 in correctness, groundedness, completeness, relevance, and fluency. By making expert knowledge actionable and accessible, SafeMate has the potential to improve public readiness, accelerate decision-making, and ultimately save lives during critical events.
@@ 1 Introduction
In light of the above, we introduce SafeMate, a retrieval-augmented, multimodal AI assistant designed to enhance both emergency preparedness and response. Unlike conventional EDSS or static checklists, SafeMate delivers adaptive, verified, and structured outputs using advanced reasoning and knowledge retrieval methods.
SafeMate is built on a modular agent architecture enabled by the Model Context Protocol (MCP)(Anthropic, 2024). MCP provides a framework for orchestrating multiple external tools and services, such as document retrievers, geospatial databases, public preparedness knowledge bases, and search APIs, under a shared semantic interface. This architecture allows the agent to dynamically route queries to the appropriate tool, retrieve domain-specific information, and generate a unified, context-sensitive response. Unlike monolithic pipelines, this modularity supports rapid integration of new knowledge sources and tools from diverse domains, ranging from legal preparedness guidelines to localized infrastructure plans.
@@ 2.4 Cross-Domain Applications of Large Language Models
Across these domains, one clear insight emerges: LLM-based systems achieve greater reliability and impact when coupled with structured reasoning, external verification, and modular extensibility, which are core tenets embodied by SafeMate’s design.
@@ 3.1 System Architecture Overview
As illustrated in Figure 1, the SafeMate system is organized into a modular architecture centered around an LLM and orchestrated via the Model Context Protocol (MCP). The MCP Client layer, implemented using LangChain and LangGraph, manages interaction between the language model and external tools. It invokes FAISS-based retrieval over an embedded knowledge base for static information and routes dynamic tool calls, such as YouTube search, Google Maps rendering, or weather APIs, through LangChain MCP adapters. These tools are accessed via the MCP Server infrastructure, such as Smithery. A FastAPI backend exposes the LLM agent to users and facilitates requests via both web and geospatial interfaces. The response pipeline culminates in the LLM composing a multimodal, context-aware output, grounded in retrieved documents and enriched with tool-based augmentations.
@@ 3.3 Low-Hallucination Language Model: o3-mini
In emergency response applications, the accuracy of information is paramount. To minimize the risk of generating incorrect or misleading content, we selected OpenAI’s o3-mini language model for SafeMate. According to evaluations using Vectara’s Hughes Hallucination Evaluation Model (Bao et al., 2024), o3-mini exhibited a hallucination rate of 0.8%, outperforming other models such as 4o (1.4%) and DeepSeek r1 (14.3%).
@@ 3.6 Hierarchical Retrieval with RAPTOR
To effectively manage and retrieve information from extensive and complex emergency documents, SafeMate integrates the RAPTOR framework. RAPTOR constructs a hierarchical tree structure through recursive embedding, clustering, and summarization of text chunks, enabling multi-level abstraction and efficient information retrieval. The processing detail of RAPTOR is presented in Figure2.
@@ 3.6.2 Dimensionality Reduction and Clustering
$p_{ij}=\exp\left(-\frac{\|\mathbf{x}_{i}-\mathbf{x}_{j}\|^{2}}{\sigma_{i}}\right)$ (2)
@@
$q_{ij}=\left(1+a\|\mathbf{y}_{i}-\mathbf{y}_{j}\|^{2b}\right)^{-1}$ (3)
@@
$\mathcal{L}_{\text{UMAP}}=\sum_{i\neq j}\left(p_{ij}\log\frac{p_{ij}}{q_{ij}}+(1-p_{ij})\log\frac{1-p_{ij}}{1-q_{ij}}\right)$ (4)
@@
$p(\mathbf{x}_{i})=\sum_{k=1}^{K}\pi_{k}\,\mathcal{N}(\mathbf{x}_{i}\mid\boldsymbol{\mu}_{k},\boldsymbol{\Sigma}_{k})$ (5)
@@
$\text{BIC}=\ln(N)\cdot k-2\cdot\ln(\hat{L})$ (6)
@@ 4.1 Evaluation and Verification
Figure 3 illustrates our evaluation framework, where GPT serves as an automated evaluator to assess the quality of SafeMate’s generated responses(Liu et al., 2023). Given an emergency context that includes structured meta-information (e.g., disaster type, time, location) and supporting documents, a generator module synthesizes a question and a corresponding answer. SafeMate then generates its own response to the same question. Both responses are passed to an LLM-based evaluator (e.g., GPT-4), which scores the outputs across five dimensions: correctness, groundedness, completeness, relevance, and fluency. This evaluation loop enables scalable, consistent, and multi-criteria assessment of SafeMate’s response quality, particularly in settings where human evaluation is costly or infeasible.
@@ 4.2 Analysis
Groundedness and Hallucination Resistance
@@ 5 Limitations and Future Work
SafeMate has primarily been tested using U.S.-centric emergency guidelines (FEMA, CDC). This restricts its applicability to other regions unless localized knowledge bases are incorporated. In future work, we aim to support multilingual knowledge bases and culturally adaptive responses. In addition, although retrieval accuracy and grounding have been evaluated, large-scale user testing in real disaster drills or field conditions remains future work. We plan to conduct scenario-based evaluations in collaboration with municipal emergency offices to quantify impact on decision-making speed and public trust.
@@ 6 Ethical Considerations
First, while SafeMate incorporates hallucination mitigation strategies, including the use of the o3-mini model and double-checking with retrieved evidence, no generative model is entirely free from factual errors. Misinterpretations of ambiguous queries, reliance on outdated documents, or edge-case generation failures could lead to inappropriate recommendations. To minimize potential harm, all outputs are explicitly labeled as informational and not legally or medically prescriptive.
@@ 7 Conclusion
We presented SafeMate, a retrieval-augmented, tool-integrated AI assistant designed to make emergency preparedness and response information more accessible, reliable, and actionable for the general public. Addressing the limitations of traditional emergency systems, namely their reliance on static documents, professional knowledge, and poor real-time usability, SafeMatete integrates modular components including the Model Context Protocol (MCP), FAISS-based similarity retrieval, RAPTOR-style hierarchical summarization, and low-hallucination LLMs.
@@
Data Availability
Data will be made available on request.